OpenStudy (anonymous):

Help! How do you solve this? Find g(a+1) when g(x)=5x+1.

OpenStudy (alexwee123):

substitute a+1 for x into f(x)

OpenStudy (helder_edwin):

Remember when you have a function g(x) the x a "mute" letter. Whatever you write between the parentheses should go where the x are.

OpenStudy (alexwee123):

so g(a+1)=5(a+1)+1 =5a+6

OpenStudy (anonymous):

Thank you!